# HeadCraft: Modeling High-Detail Shape Variations for Animated 3DMMs

### v1

@@ Abstract
Current advances in human head modeling allow to generate plausible-looking 3D head models via neural representations. Nevertheless, constructing complete high-fidelity head models with explicitly controlled animation remains an issue. Furthermore, completing the head geometry based on a partial observation, e.g. coming from a depth sensor, while preserving details is often problematic for the existing methods. We introduce a generative model for detailed 3D head meshes on top of an articulated 3DMM which allows explicit animation and high-detail preservation at the same time. Our method is trained in two stages. First, we register a parametric head model with vertex displacements to each mesh of the recently introduced NPHM dataset of accurate 3D head scans. The estimated displacements are baked into a hand-crafted UV layout. Second, we train a StyleGAN model in order to generalize over the UV maps of displacements. The decomposition of the parametric model and high-quality vertex displacements allows us to animate the model and modify it semantically. We demonstrate the results of unconditional generation and fitting to the full or partial observation.
Figure 1: We present HeadCraft [^2] Project page is available at https://seva100.github.io/headcraft., a generative model for highly-detailed human heads, ready for animation. Our method is trained on 2D displacement maps collected by registering a parametric template head with free surface displacements to a large set of 3D head scans. The resulting model is highly versatile, which can be demonstrated by fitting the latent code of the model to an arbitrary depth observation.
@@ 1 Introduction
The advent of neural representations, such as NeRFs (Neural Radiance Fields) [38] and SDFs (Signed Distance Functions) [40], has revolutionized the field of 3D modeling by enabling the generation of remarkably realistic 3D head models. These models have found their applications in various domains, including computer graphics, virtual reality, and digital entertainment. The ability to create lifelike 3D head models is crucial for many applications, ranging from video game character design to virtual try-on experiences and medical simulations. While contemporary techniques have made significant strides in achieving high-quality 3D head models, a series of challenges persist, impeding the seamless integration of these models into real-world applications. In particular, one of the primary challenges in the realm of 3D head modeling lies in constructing a neural representation amenable to animation and tracking, while preserving enough detail.
In particular, recently introduced implicit generative models for 3D, such as pi-GAN [14], EG3D [15], or StyleSDF [39], fit the distribution of human faces with extraordinary detail, and e.g. in [7, 16], the modeled region is extended to the whole head. To introduce the animation capabilities, these methods require ad-hoc modifications such as introducing deformations to canonical space [8] or driving the rendering by a semantic face mask [49, 56]. The methods mostly focused on both modeling and controlling the geometry are typically SDF-based and require learning a separate latent space for expressions [60], oftentimes with the aid of deformations [26, 61]. At the same time, approaches that fit a neural representation to a ”talking head”-style video, enable explicit articulation by introducing the parametric model guidance [27, 64, 65].
@@
Inspired by the combination of these ideas, in this research, we introduce a generative model that allows for animation and tracking and preserves high level of detail. At the heart of our approach lies the idea of combining an explicit parametric head model (FLAME [36]) with surface displacements complementing the low geometry detail of the head model. FLAME is an example of a 3D Morphable Model [10, 20] with a fixed set of vertices and fixed topology, constructed as a linear statistical model over the heads with point-to-point correspondence and further controlled by shape and expression latent codes. To obtain the necessary training data, we register a highly subdivided FLAME mesh template with free vertex displacements to all 3D head scans in the NPHM [26] dataset. To facilitate as high level of detail in the displacements as possible, they are fitted in two steps. First, the optimization problem is solved for vector displacements with strong regularization that penalizes very hard for self-intersections of the deformed mesh regions. Afterwards, a separate optimization step refines the displacements only along the normals of the deformed vertices, while keeping the regularization weight low. These displacements are baked into a predefined UV layout. Finally, we train a StyleGAN2 [31] model to generalize over this set of baked 2D displacement maps. This novel architecture allows us to operate at a resolution higher than the conventional FLAME template, enabling the generation of highly detailed and animatable 3D head models.
@@
To validate the efficacy and practical utility of our approach, we evaluate it in several settings. The diversity and fidelity of the generated 3D head meshes is quantitatively and visually compared to other methods w.r.t. the real head scans from the FaceVerse dataset [55], both in UV space and rendered image space. We also explore the applicability of our approach in fitting the latent representation of the generative model to complete or incomplete point cloud data and demonstrate its animation and manipulation capabilities.
@@
Our contributions are as follows:
  • We introduce a two-stage registration procedure to craft high-detail displacement maps on top of 3DMMs from 3D scanning data.
  • We propose a generative model over displacement maps to not only enhance the low-frequency geometry of FLAME with details but also extend its shape space to all kinds of hairstyle variations.
  • We demonstrate the versatility of our method through unconditional sampling, interpolation, semantic hair transfer, and conditioning by depth map or a complete scan.
@@ 2 Related Work
Many recent solutions to computer vision problems involving human bodies are built on statistical body models. They are the foundation for building personalized avatars [27, 66, 4, 25, 5], motion tracking [51, 22], scan registration [26], controlling image synthesis [50], and many more. Their line of research divides into two major branches.
@@
Mesh-based Models. Pioneering work in the field [10] proposed 3D morphable models (3DMMs) for identity, expression, and appearance representation of human faces. Their model is built around a 3D template mesh and linear parametric blendshapes derived from PCAs over scan data. With new datasets and registration procedures, their work has been extended from faces to heads [36, 35, 42], hands [46], full-bodies [37], or combinations of these [41, 57]. The template mesh has a fixed topology which provides consistent UV unwrapping and enables fitting to know surface correspondences, e.g. semantic regions and landmarks. Yet, it limits the representative power w.r.t. the overall shape and the level of detail beyond what is provided by the template. Downstream approaches compensate this by optimizing displacements [32, 27, 4, 12, 5, 59, 35] or additional implicit geometry [23, 13] on top of the mesh. Displacements are applied either per-vertex individually [32, 27, 4, 5] or as a displacement map over the whole surface using the consistent UV unwrapping of the template [59, 35]. Our method follows this idea by learning a generative model over displacements maps while exploiting the animation model and surface correspondences of the underlying 3DMM. We demonstrate that learning displacements maps on top of a highly-subdivided template allows to model fine details and, compared to previous works, to introduce significant shape variations.
@@
Implicit Models. The recent success of implicit SDFs [40] and neural radiance fields [38] in 3D modeling has also motivated applying them for statistical body models. Most implicit models learn shape and appearance in a canonical reference space [26, 60, 6, 29, 43] or as displacements on top of an existing model [61]. For better generalization and detail preservation, some approaches use a composition of implicit SDFs to model the canonical space [6, 26]. Articulation and animation is modelled either directly in canonical coordinates [29, 61], through implicit deformation fields [26, 60], explicit joints [6] or blendshape deformations borrowed from explicit 3DMMs [64, 56]. While the aforementioned methods rely on multi-view data and aligned 3D scans, a separate line of research demonstrates that statistical shape and appearance priors can also be learned from unstructured image collections [14, 15, 45, 39, 67].
@@
Implicit approaches do not rely on topology and shape templates. This allows them to fit more detail and complex shapes such as hair [26, 61, 60] and even glasses [15, 39]. Yet, it prevents consistent surface correspondences between different samples which need to be explicitly learned [61, 6]. As our approach uses a mesh-based template, it does not suffer from these limitations and has an explicit model for animation. Still, we are able to show that we can provide a comparable level of detail as implicit methods by learning high-resolution displacement maps on top of a highly subdivided template topology.
@@
Figure 3: Visual comparison of fidelity and diversity of the meshes generated by various methods. For Ours, random FLAMEs are sampled from Gaussian distribution with statistics calculated over the NPHM dataset; same for the PCA baseline pre-fitted to our UV registrations. Meshes from NPHM are obtained by sampling the latent codes and running marching cubes over the generated SDF representations. We demonstrate higher variability of produced head geometry and better details than the other methods. Electronic zoom-in recommended.
3 Method
We describe the stages of our method in several parts. Subsec. 3.1 describes the registration procedure for FLAME with surface displacements. Subsec. 3.2 describes learning the generative model. The overview of the approach is presented in Fig. 2.
@@ 3.1 Displacements registration procedure
The purpose of this step is to align the FLAME [36] head template with displacements to each scan of the 3D human head dataset. Let us consider a single scanned mesh $\mathcal{P}=(V^{\textrm{gt}},\mathcal{F}^{\textrm{gt}})$ with vertices $V^{\textrm{gt}}\in\mathrm{R}^{|V^{\textrm{gt}}|\times 3},$ and faces $\mathcal{F^{\textrm{gt}}}\in\mathrm{R}^{|\mathcal{F}^{\textrm{gt}}|\times 3}$.
@@ 3.2 Generative model
The described registration procedure allows us to relax the problem of 3D head geometry generation into a problem of generation of 2D UV displacement maps, which allows us to apply a 2D generative model. We have selected StyleGAN2 for that purpose due to its capability of generalizing over relatively small datasets of images [30, 63, 1] while maintaining close-to-SoTA image generation capabilities [31]. The model consists of a mapping network and a generator network, which we will refer to as $f(z)$ together, where $z\in\mathcal{Z}\subset\mathbb{R}^{\textrm{D}}$ is a latent code sampled from a standard normal distribution during training. The generator produces a UV offset map $U=f(z)$, which we can apply to an arbitrary (anyhow densely subdivided) FLAME template $F=(V,\mathcal{F},\mathcal{U}_{\mathcal{F}})$ by querying the map $U$ with its texture coordinates $\mathcal{U}_{\mathcal{F}}$ to obtain the respective vertex displacements. We later demonstrate visually that the generated displacements could be applied to an arbitrary FLAME template.
[Uncaptioned image] [Uncaptioned image]
Figure 4: Comparison of the FLAME layouts. The standard, commonly used unwrapping for FLAME (left) features a seam corresponding to the vertical line in the back of the head and pays more attention to the facial region than to the scalp. In the hand-crafted custom layout that we employ (right), a different seam around the face border is selected, thus making the regions of face and scalp separated and of similar size, which simplifies modeling complex geometry such as long hair without breaks.
Post-processing. Since the UV map $U$ is generated in the UV layout that contains a seam, we expect StyleGAN to resolve it in general, i.e. produce similar displacements in the face and scalp region near the same seam vertex. Still, there is no dedicated supervision during StyleGAN training that ensures that it always happens and that the border is preserved pixel-perfect. Because of that, before querying the UV map $U$, we ensure the smoothness of the seam by equalizing the displacements at its borders and blending the displacements in the small vicinity inside and outside of it. The sampling errors near the seam are fixed by filling in all the empty space in the map with the displacements copied from the nearest neighbor pixels in the valid regions.
@@ 4 Experiments
4.1 Training procedure
@@
Our method is trained on the private version [^3] The private version was provided to us by the authors of NPHM [26] and is going to be released to the public. The public version, available at the time of the publication, is roughly a 30% subset of the private version. of the NPHM [26] dataset, consisting of 6975 high-resolution scans of 327 diverse identities captured by two 3D scanners. For the registration procedure, we use Adam optimizer with learning rate of $3\cdot 10^{-2}$ for the first stage and $3\cdot 10^{-4}$ for the second stage. The hyperparameters $\boldsymbol{\lambda}_{\textrm{Stage 1}}=(\lambda_{\textrm{Stage 1}}^{\textrm{Chamfer}},\lambda_{\textrm{Stage 1}}^{\textrm{edge}},\lambda_{\textrm{Stage 1}}^{\textrm{lapl}})$ equal to $(2\cdot 10^{3},2\cdot 10^{5},10^{4})$. For the second stage, $\boldsymbol{\lambda}_{\textrm{Stage 2}}=(2\cdot 10^{4},2\cdot 10^{4},10^{4})$. In the Chamfer loss, we additionally apply correspondences pruning by distance of 1.0, which defines that all the correspondences between source and target with the distance more than 1.0 in the NPHM coordinate system are automatically discarded. This has been introduced for more consistent gradual learning of displacements, such that at each optimization step, only the nearest points affect the deformation learning.
@@
For the generative model training, we use StyleGAN2 implementation with all augmentations turned off (since they wouldn’t yield valid UV maps in our case), 8 mapping network layers and a high gradient penalty of 4.0. The learning rates are $2\cdot 10^{-3}$ for the generator and $1\cdot 10^{-3}$ for the discriminator. We train it for 72K steps with the batch size of 8 and the resolution of the UV maps of $256\times 256$.

### v2

@@ Abstract
Current advances in human head modeling allow to generate plausible-looking 3D head models via neural representations, such as NeRFs and SDFs. Nevertheless, constructing complete high-fidelity head models with explicitly controlled animation remains an issue. Furthermore, completing the head geometry based on a partial observation, e.g. coming from a depth sensor, while preserving a high level of detail is often problematic for the existing methods. We introduce a generative model for detailed 3D head meshes on top of an articulated 3DMM which allows explicit animation and high-detail preservation at the same time. Our method is trained in two stages. First, we register a parametric head model with vertex displacements to each mesh of the recently introduced NPHM dataset of accurate 3D head scans. The estimated displacements are baked into a hand-crafted UV layout. Second, we train a StyleGAN model in order to generalize over the UV maps of displacements, which we later refer to HeadCraft. The decomposition of the parametric model and high-quality vertex displacements allows us to animate the model and modify the regions semantically. We demonstrate the results of unconditional sampling, fitting to a scan and editing. The code and data are available at https://seva100.github.io/headcraft.
Figure 1: We present HeadCraft, a generative model for highly-detailed human heads, ready for animation. Our method is trained on 2D displacement maps collected by registering a parametric template head with free surface displacements to a large set of 3D head scans. The resulting model is highly versatile and its latent code can be fit to an arbitrary depth observation.
@@ 1 Introduction
The ability to create lifelike 3D head models is crucial for many applications, ranging from video game character design to virtual try-on experiences and medical simulations. 3D Morphable Models (3DMMs) [6] constitute an essential and robust tool for basic 3D head geometry estimation and tracking that enables its further reconstruction and animation. Furthermore, 3DMMs, along with their consistent UV parameterization of the surface, are commonly used in constructing virtual avatars as an approach for approximate surface representation or regularization  [37, 26, 4, 66].
Constructing (neural) 3DMMs capable of representing the diverse distribution over human heads while disentangling identity and expressions and capturing a high degree of detail is challenging. At the same time, such neural representation would ideally be compatible with well-established methods for animation and tracking, which are usually mesh-based and are required to fulfill certain real-time constraints.
Despite the recent progress of implicit representations, such as neural radiance fields (NeRFs) [40] and signed distance functions (SDFs) [42], the most prominent 3DMMs [6, 36, 44] are based on a template mesh (i.e. feature explicit geometry), and principal component analysis (PCA), to represent identity and expression variations. Mesh-based 3DMMs can usually be robustly fitted to videos, easily animated and integrate well with established graphics pipelines. However, these approaches are fundamentally limited by both the mesh resolution and representational capacity of its underlying (multi-)linear statistical model. In our work, we improve on both of these aspects and leverage recent advances in generative image modeling by using StyleGAN2 [31] to predict highly detailed geometry in the UV space, which is independent of the mesh resolution.
A different line of work attempts to build 3DMMs based on neural SDFs instead of meshes (e.g.  [61, 24]), thereby enabling reconstruction at arbitrary resolutions. However, the incorporation of such implicit representations is limited due to a lack of compatibility with standard graphics systems and animation tools. Furthermore, SDF-based approaches often require costly evaluations, e.g. via marching cubes [39], to extract an implicit surface, which can hinder the real-time application of these methods.
@@
Inspired by the combination of these ideas, in this research, we introduce a generative model that allows for animation and tracking and preserves a high level of detail. At the heart of our approach lies the idea of combining an explicit parametric head model (FLAME [36]) with surface displacements complementing the low geometry detail of the head model. FLAME is an example of a 3D Morphable Model [6, 19] with a fixed set of vertices and fixed topology, constructed as a linear statistical model over the heads with point-to-point correspondence and further controlled by shape and expression latent codes. We register a highly subdivided FLAME mesh template with free vertex displacements to all 3D head scans in the NPHM [24] dataset to obtain the necessary training data. To facilitate as high level of detail in the displacements as possible, they are fitted in two steps. First, the optimization problem is solved for vector displacements with strong regularization that penalizes very hard for self-intersections of the deformed mesh regions. Afterwards, a separate optimization step refines the displacements only along the normals of the deformed vertices while keeping the regularization weight low. These displacements are baked into a predefined UV layout. Finally, we train a StyleGAN2 [31] model to generalize over this set of baked 2D displacement maps. This novel architecture allows us to operate at a resolution higher than the conventional FLAME template, enabling the generation of highly detailed and animatable 3D head models.
@@
To validate the efficacy and practical utility of our approach, we evaluate it in several settings. The diversity and fidelity of the generated 3D head meshes are quantitatively and visually compared to other methods w.r.t. the real head scans from the FaceVerse dataset [56], both in UV space and rendered image space. We also explore the applicability of our approach in fitting the latent representation of the generative model to complete or incomplete point cloud data and demonstrate its animation and manipulation capabilities.
@@
To summarize, our contributions are as follows:
  • We introduce a two-stage registration procedure to craft high-detail displacement maps on top of 3DMMs from 3D scanning data. This enables the application of 2D generative models to tackle the 3D generative task.
  • We propose a generative model operating in the displacement maps domain to enhance the low-frequency geometry of FLAME with details and extend its shape space to a range of variations.
  • We demonstrate the versatility of our method through unconditional sampling, interpolation, semantic geometry transfer, and 3D completion based on partial depth observations.
@@ 2 Related Work
Many recent solutions to computer vision problems involving human bodies and heads are built on statistical body models, forming the foundation for building personalized avatars [26, 66, 1, 23, 2], motion tracking [52, 21], scan registration [24], controlling image synthesis [51], and more. Their line of research is divided into two major branches.
@@
Mesh-based Models. Pioneering work in the field [6] proposed 3D morphable models (3DMMs) for the human faces’ identity, expression, and appearance representation. Their model is built around a 3D template mesh and linear parametric blendshapes derived from PCAs over 3D scan data. With new datasets and registration procedures, their work has been extended from faces to heads [36, 35, 46], hands [50], full bodies [38], or combinations of these [43, 58]. The template mesh has a fixed topology which provides consistent UV unwrapping and enables fitting to know surface correspondences, e.g. semantic regions and landmarks. Yet, it limits the representative power w.r.t. the overall shape and level of detail beyond what the template provides. Downstream approaches compensate this by optimizing displacements [32, 26, 1, 9, 2, 60, 35] or additional implicit geometry [22, 10] on top of the mesh. Displacements are applied either per-vertex individually [32, 26, 1, 2] or as a displacement map over the whole surface using the consistent UV unwrapping of the template [60, 35, 21]. Some approaches infer the displacement maps from images or texture reprojections for specific individuals [60, 21]. By making use of high-quality scans, the authors of [35] demonstrate that GANs [25, 29, 31] for image generation can be leveraged to learn a generative model over displacement maps in the face region. Our method takes this idea further by learning a generative model for displacements over the whole head. It showcases that even the surface geometry of large and complicated hairstyles can be represented with high fidelity. The resulting outputs of our full-head generative model provide quality approaching the scan level while exploiting the UV surface correspondences of the underlying 3DMM and enabling further animation through the rigging of the template.
@@
Implicit Models. The recent success of implicit SDFs [42] and neural radiance fields [40] in 3D modeling has also motivated applying them for statistical body models. Most implicit models learn shape and appearance in a canonical reference space [24, 61, 3, 28, 47] or as displacements on top of an existing model [62]. For better generalization and detail preservation, some approaches use a composition of implicit SDFs to model the canonical space [3, 24]. Articulation and animation are modeled either directly in canonical coordinates [28, 62], through implicit deformation fields [24, 61], explicit joints [3] or blendshape deformations borrowed from explicit 3DMMs [65, 57]. While the aforementioned methods rely on multi-view data and aligned 3D scans, a separate line of research demonstrates that statistical shape and appearance priors can also be learned from unstructured image collections [11, 12, 49, 41, 17].
@@
Implicit approaches do not rely on topology and shape templates. This allows them to fit more detail and complex shapes such as hair [24, 62, 61] and even glasses [12, 41]. Yet, it prevents consistent surface correspondences between different samples, which need to be explicitly learned [62, 3]. As our approach uses a mesh-based template, it does not suffer from these limitations and has an explicit model for animation. Still, we are able to show that we can provide a comparable level of detail as implicit methods and also model hair surface geometry which has not been achieved with a generative, explicit shape model before.
@@
Figure 3: Visual comparison of fidelity and diversity of the meshes generated by various methods. For Ours, random FLAMEs are sampled from Gaussian distribution with statistics calculated over the NPHM dataset; same for the PCA baseline pre-fitted to our UV registrations. Meshes from NPHM are obtained by sampling the latent codes and running marching cubes over the generated SDF representations. We demonstrate higher variability of produced head geometry and better details than the other methods.
3 HeadCraft
We cast the task of learning high-fidelity distribution over 3D heads as a 2D generative problem by leveraging the UV-space of an existing 3DMM [36]. This allows us to rely on the well-explored body of research on CNN-based 2D generative models [31, 30].
To this end, we register a dataset of high-end 3D head scans [24] in the FLAME head model [36] topology and bake the details into a displacement map in UV space, as explained further in Sec. 3.1. Subsequently, in Sec. 3.2, we train a 2D generative model to learn the distribution over the displacement maps, which can be converted into detailed 3D displacements compatible with the underlying FLAME model. An overview of HeadCraft is presented in Fig. 2.
@@ 3.1 Displacements registration procedure
The purpose of this step is to create a dataset of 2D displacement maps representing the geometric information beyond the representational capacity of classical 3DMMs. To this end, we compute displacements from a FLAME [36] mesh to the high quality 3D head scans from the NPHM dataset [24]. Let us consider a scanned mesh $\mathcal{P}=(V^{\textrm{gt}},\mathcal{F}^{\textrm{gt}})$ with vertices $V^{\textrm{gt}}\in\mathbb{R}^{|V^{\textrm{gt}}|\times 3},$ and faces $\mathcal{F^{\textrm{gt}}}\in\{1,\dots,|V^{\textrm{gt}}|\}^{|\mathcal{F}^{\textrm{gt}}|\times 3}$.
@@ 3.2 Generative model
The described registration procedure allows us to relax the problem of 3D head geometry generation into a problem of generation of 2D UV displacement maps, which allows us to apply a 2D generative model. We have selected StyleGAN2 for that purpose due to its capability of generalizing over relatively small datasets of images [30, 64, 45] while maintaining close-to-SoTA image generation capabilities [31]. The model consists of a mapping network and a generator network, which we will refer to as $f(z)$ together, where $z\in\mathcal{Z}\subset\mathbb{R}^{\textrm{D}}$ is a latent code sampled from a standard normal distribution during training (with truncation trick [7] at the inference time). The generator produces a UV displacement map $U=f(z)$, which we can apply to an arbitrary (anyhow densely subdivided) FLAME template $F=(V,\mathcal{F},\mathcal{U}_{\mathcal{F}})$ by querying the map $U$ with its texture coordinates $\mathcal{U}_{\mathcal{F}}$ to obtain the respective vertex displacements. The final mesh $M=(V+D(U),\mathcal{F},\mathcal{U}_{\mathcal{F}})$ is different from the subdivided FLAME only in terms of the vertex locations. We later demonstrate visually that the generated displacements could be applied to an arbitrary template.
UV layout. Importantly, we modify the UV embedding of the FLAME template mesh into a 2D plane (see the Supplementary for the illustration). Doing so results in a more favorable layout for the 2D generative model with a limited receptive field. As we demonstrate further, the large distance in the UV space of the head’s back left and right parts leads to inconsistent generations. The new layout allows us to stack the left- and right-hand sides into a single-channel displacement map with better seam alignment.
Post-processing. Since the UV map $U$ is generated in the UV layout that contains a seam, we expect StyleGAN to resolve it in general, i.e. produce similar displacements in the face and scalp region near the same seam vertex. Still, there is no dedicated supervision during StyleGAN training that ensures that it always happens and that the border is preserved pixel-perfect. Because of that, we apply Laplacian smoothing [55] to the mesh $M$ in the $K$-vertex vicinity of the seam. Additional smoothness of the face region is achieved by applying Laplacian smoothing to the facial skin, neck, scalp, eyeballs, and inner mouth region with different strength to the subdivided FLAME template in advance, before adding the displacements. Technical details are provided in the Supplementary Material.
FID ↓ KID ↓ MMD ↓ JSD ↓ COV ↑
Ours 68.00 0.065 6.51 21.33 53.85%
PCA 126.31 0.165 10.47 20.16 23.08%
NPHM 139.82 0.170 7.80 19.06 46.15%
ROME 169.65 0.204 10.02 23.19 32.69%
FLAME 198.85 0.262 12.95 23.89 5.77%
Table 1: The comparison of quality and diversity of random samples generated by each of the methods. FID and KID measure the similarity of the generated mesh renderings vs. the renderings of the ground truth meshes in FaceVerse dataset, while 3D metrics MMD, JSD, COV assess the similarity of generated and real point clouds distributions. MMD is multiplied by $10^{3}$ and JSD by $10^{2}$.
@@ 4 Experiments
We evaluate HeadCraft’s generative capabilities by examining its unconditional generation performance in Sec. 4.4 and ablate over several important aspects. Additionally, in Sec. 4.5, we demonstrate the way HeadCraft can be beneficial in several important downstream applications, such as 3D head completion from a partial point cloud, its seamless integration with the FLAME expression space, and semantic geometry transfer. Before presenting these results, we describe implementation details, as well as our chosen metrics and baselines in Sec. 4.1, 4.2, and 4.3, respectively.
4.1 Implementation details
Datasets. Our method is trained on the 3D real scans of human heads from NPHM [24] dataset, namely, 6975 high-resolution scans of 327 diverse identities captured by two 3D scanners each. To quantitatively evaluate the generative performance of all compared methods, we use FaceVerse [56] dataset as a source of the ground truth scans not intersecting with our training data by identities.
@@ 4.1 Implementation details
Registration procedure. We use Adam optimizer with learning rate of $3\cdot 10^{-2}$ for the first stage and $3\cdot 10^{-4}$ for the second stage. The hyperparameters $\boldsymbol{\lambda}_{\textrm{Stage 1}}=(\lambda_{\textrm{Stage 1}}^{\textrm{Chamfer}},\,$ $\lambda_{\textrm{Stage 1}}^{\textrm{edge}},\,\lambda_{\textrm{Stage 1}}^{\textrm{lapl}})$ equal to $(2\cdot 10^{3},2\cdot 10^{5},10^{4})$. For the second stage, $\boldsymbol{\lambda}_{\textrm{Stage 2}}=(2\cdot 10^{4},2\cdot 10^{4},10^{4})$. In the Chamfer loss, we additionally apply correspondences pruning by distance of 1.0, which defines that all the correspondences between source and target with the distance more than 1.0 in the NPHM coordinate system are automatically discarded. This has been introduced for more consistent gradual learning of displacements, such that at each optimization step, only the nearest points affect the deformation learning.
@@
Generative modeling. As the 2D generative model, we choose StyleGAN2-ADA [30] with all augmentations turned off (since they wouldn’t yield valid UV maps in our case) and 8 mapping network layers. To stablize the GAN training, we utilize a high gradient penalty of 4.0 for the discriminator. The learning rates are $2\cdot 10^{-3}$ for the generator and $1\cdot 10^{-3}$ for the discriminator. We train it for 95K steps with the batch size of 8 and $256\times 256$ image resolution.
4.2 Metrics
To evaluate the quality of the generated samples in section 4.4, we rely on both 2D metrics computed on renderings, as well as, on metrics directly compute in 3D. In the following we describe all employed metrics in details.
Firstly, to evaluate the visual plausibility of the generated geometry, we render 2195 ground truth meshes from FaceVerse and the same number of meshes generated by each method with highly metallic material from eight distinct viewpoints, uniformly sampled along the circular trajectory in the horizontal plane. The FID [27] and KID [5] perceptual metrics are calculated for all generated and ground truth renderings from a given viewpoint and then macro-averaged over eight viewpoints.
Secondly, we compare the distributions of point clouds sampled from the generated and ground truth meshes. To do that, we sample 10K points from each of the 2195 generated and the same number of ground truth meshes and calculate several 3D similarity metrics. Jensen-Shannon Divergence (JSD) is evaluated by comparing the distributions of generated and ground truth points, splat into a voxel grid (in our case, of $512^{3}$ voxels). Minimum Matching Distance (MMD) is a measure of 3D object realism that, for each ground truth sample, involves evaluating the distance to the most similar sample in the generated set. Similarly, Coverage (COV) indicates the percentage of the ground truth samples, for which the nearest neighbor among all ground truth and generated samples falls into the generated set. More detailed description of the 3D metrics can be found in [59, 20].
4.3 Baselines
We compare against the recent state-of-the-art head modelNPHM [24], which uses neural SDFs to represent the head geometry. We additionally compare againstROME [32], an alternative approach that models complete head geometry including hair using a FLAME template mesh. ROME is trained in an unsupervised fashion on the large-scale video dataset VoxCeleb2 [13].
Furthermore, we compare against a PCA-based baseline, whereas a linear PCA basis is fitted to our UV displacement maps, and provide the numbers for random FLAME samples without added displacements as a reference.
While NPHM, PCA baseline, and Ours have been fitted to exactly the same training dataset, for ROME, the checkpoint from the public repository has been used.